# Robust variational neural posterior estimation for simulation-based inference

### v1

@@ 2.3 Misspecification in SBI
The simulator, $p(\bm{x}_{\rm sim}|\,\bm{\theta})$, is said to be misspecified if the true data-generating process does not fall within the family of distributions defined by the simulator on the support of the prior of $\bm{\theta}$. That is, $q^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\rm{supp}(p(\bm{\theta}))\}$ (Cannon et al., 2022). For single observations, the definition of misspecification was extended to using summary statistics (Kelly et al., 2024) by defining $\bm{b}(\bm{\theta})=\mathbb{E}_{p(\bm{x}_{\rm sim}\mid\bm{\theta})}[\iota_{\omega}(\bm{x}_{\rm sim})]$ and $\bm{b}_{j}=\mathbb{E}_{q^{*}(\bm{x}_{\rm obs})}[\iota_{\omega}(\bm{x}_{\rm obs}^{(j)})]$, for each $\bm{x}_{\rm obs}^{(j)}\in O$ where $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ is the set of observations each associated to a different, unknown true $\bm{\theta^{(j)}}$ value. Then the the simulator is misspecified if there is no $\{\bm{\theta}\}_{i=1}^{N_{\rm obs}}$ in the support of the prior for which $\bm{b}(\bm{\theta}^{(j)})=\bm{b}_{j}$ for each $j$.
In this paper, we adopt the alternate, but similar, definition of misspecification provided in Wehenkel et al. (2025), which is more aligned with amortised SBI. They define a simulator to be misspecified if $\exists\mathcal{S}\subseteq\Theta\times\mathcal{X}_{\rm obs}:\forall(\bm{\theta},\bm{x}_{\rm obs})\in\mathcal{S}$
@@
This definition aligns with the amortised SBI task as it identifies misspecification as a set-wise phenomenon. Our goal is to recover robust and reliable posterior inference for all $\bm{x}_{\rm obs}\in\mathcal{X}_{\rm obs}$. This definition does, however, ignore situations when the prior distribution is misspecified. Our method can handle prior misspecification more favourably through traditional Bayesian model iterations, provided the support of the proposed prior is a subset of the support of the original prior.
@@ 2.4 Importance weighted autoencoders
Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, Kingma & Welling 2013), which introduces a latent space $\mathcal{Z}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\xi}}(\bm{x}|\,\bm{z})$ and $p_{\bm{\phi}}(\bm{z}|\,\bm{x})$, known as the decoder and the encoder respectively. The vanilla VAE objective aims to maximise the evidence of the data via the evidence lower bound
$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{p_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\bm{\xi}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\!\left(p_{\bm{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (5)
with respect to $\bm{\xi}$ and $\bm{\phi}$, where $p(\bm{z})$ is a prior over the latent parameter. However, using this bound has been shown to induce mode-seeking behaviour, leading to overly simplified representations and poor inference. Burda et al. (2015) introduced IWAE, which is based on a strictly tighter evidence lower bound derived from importance sampling. The log-evidence for a single point is given as
$\log p_{\bm{\theta}}(\textbf{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\xi}}\left(\textbf{x},\textbf{z}^{(l)}\right)}{p_{\bm{\phi}}\left(\textbf{z}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\xi},\bm{\phi};\textbf{x}\right)$ (6)
$\text{with}\quad\textbf{z}^{(l)}\sim p_{\bm{\phi}}\left(\textbf{z}|\textbf{x}\right),$ (7)
which is a mass-covering objective that targets the evidence.
@@ 3 Method: RVNP
Motivated by a desire to have a method that can recover robust amortized posterior inference with a fixed embedding space such that the simulation-to-reality gap is bridged with an interpretable, flexible error model, we extend the approaches of Ward et al. (2022) and Kelly et al. (2024) to amortized SBI, and use an importance weighted autoencoder (Burda et al., 2015) amortized variational inference scheme to define robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T). In what follows, we drop the explicit embedding notation $\iota_{\omega}$ unless necessary. Our goal is to use a normalizing flow parametrized by $\bm{\phi}$, $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$, that can approximate the true posterior distribution. Algorithm 1 provides an overview of the RVNP and RVNP-T algorithms. We chose a fixed simulation budget of $N_{\rm sim}=100,000$ for each task, with $10\%$ retained for validation.
RVNP builds on variational methods for solving the inverse problem under a learnt likelihood in SBI (Glöckler et al., 2022), and relies on a conditional neural density estimator to approximate the likelihood from a fixed budget of simulations from the simulator model. This overcomes the computational expense of using Hamiltonian Monte Carlo Neal (2011) or other Markov Chain methods to jointly infer the parameters of interest and the parameters of the error model. Furthermore, variational methods allow us to use more expressive models for the error model distribution. We assume a pre-training or fixing of the neural statistic embedding $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ occurred.
@@ 3.1 Generative model
Throughout, we assume that the neural embedding or summary statistics are pre-trained and are not jointly learnt. The first step of RVNP is to train the normalising flow $p_{\bm{\Psi}}(\bm{x}|\,\bm{\theta})$ to approximate the likelihood from the generated samples from the simulator using the NLE objective (Equation 3). Once the normalizing flow has been trained, we assume that $p(\bm{x}_{\rm sim}|\,\bm{\theta})\approx p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ and include the surrogate in the forward model.
@@ 3.2 Variational posterior
$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]\,p(\bm{\theta}^{(l)})p(\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha}))}{p_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]$ (10)
where $\bm{\theta}^{(l)}\sim p_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, $\bm{\alpha}$ are the weights of the neural network $\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]$ using a Monte Carlo (MC) estimate. We use the logsumexp function to ensure the MC estimates are stable.
@@ 3.3 Posterior tuning
$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})}[\log p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (11)
This final objective can be identified with the noisy neural posterior estimation (NNPE, Ward et al. 2022) objective and the error augmentation method suggested by Cranmer et al. (2020). However, our error model does not have to be globally fixed and has been inferred using variational inference.
@@ 3.4 Error modelling
In the experiments, we adopt two error models. The default error model that we adopt in RVNP is given by the Gaussian covariance matrix
@@
a neural network that outputs the diagonal components of the covariance matrix and the non-diagonal components $\bm{\Lambda}$ are globally learnt. We also include a global error model.
@@
defined through a full rank Gaussian covariance matrix that is parametrised in terms of a Cholesky decomposition. This error model is constant across the parameter space and does not explicitly depend on $\bm{\theta}$. Our method generalises very easily to include any inductive bias that we believe explains the simulation-to-reality gap.
Our approach can be understood in terms of the simulator defining a population prior $p(\bm{\theta},\bm{x}_{\rm sim})=p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})p(\bm{\theta})$, while the error model can be identified with the likelihood of $\bm{x}_{\rm sim}$ for the observed data. In this framework, the simulator model describes our prior understanding of the true DGP. In this paper, we allow only for Gaussian error-type models that allow the forward model to inflate the noise on the output of the simulator. This allows us to account for the misspecification using a Gaussian forget mechanism.
Including an error model is an assumption about the true DGP, and the exact corruption may be unknown. However, we follow the Box principle of all models are wrong, some are useful (Box, 1976) and attempt to forward model the DGP using the simulator and an error model term that can account for corruption in the data. Error modelling in robust SBI is not new; Ward et al. (2022) introduced a spike and slab error model, which assumes that a summary statistic is either well-specified or misspecified. The parameter adjustment method of Kelly et al. (2024) can be thought of an error model $p_{\bm{\xi}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\xi})$, integrating with respect to $\bm{x}_{\rm sim}$ first.
Adjustment parameters will move the observed point to a region of high probability with respect to the simulated samples. On a point-by-point basis, there are infinitely many solutions to this problem, and there is no guarantee of where an OOD point should be mapped back to in the original sample without a high volume of data. Furthermore, the spike-and-slab error model invokes the assumption that the true corruption model is unknown, and that the misspecification forces us to forget certain summary statistics if they lie OOD. This assumption relies on the misspecification occurring along individual axes, which is unrealistic in many situations and is subject to the hyperparameter choices of the spike-and-slab error model. In general, misspecification has no guarantee of being represented exactly along the given axes, even if the components represent summary statistics. To consider why, let $p_{\gamma}(\bm{x})$ be the probability density function parametrised by $\gamma$ and assume that $T(\bm{x})$ represents a sufficient statistic of $\bm{\theta}$. By the Fisher-Neyman factorization theorem (Hogg et al. 2005, pp 376-377), there exist non-negative functions $k_{\gamma}$ and $h$ such that $p_{\gamma}(\bm{x})=h(\bm{x})k_{\gamma}(\bm{x})$. If $g$ is a bijection, then
$f_{\gamma}(x)=h(x)\,k_{\gamma}(g^{-1}(g(T(x)))),$ (14)
$f_{\gamma}(x)=h(x)\,(k_{\gamma}\circ g^{-1})(g\circ T(x)),$ (15)
and, therefore, $g\circ T(x)$ is also a sufficient statistic. In this paper, we choose to model a full covariance matrix to account for misspecification that may not occur exactly along each axis. While suitable pre-processing may help address this issue, we choose to model the full covariance and infer this.
@@ 3.5 Prior over the correction model parameters
In RVNP, a prior over the error model parameters can be introduced to regularise the influence of the error model and make the inference more robust in the single-posterior case. Adopting an error model makes the generative model overparameterized, and the inference process will be heavily dependent on the prior assumptions about misspecification in the single-point inference. In our experiments, we do not use any prior distribution over the error model parameters. This allows us to focus on the robustness imparted by multiple observations.

### v2

@@ 2.3 Misspecification in SBI
The simulator, $p(\bm{x}_{\rm sim}|\,\bm{\theta})$, is said to be misspecified if the true data generative process does not fall within the family of distributions defined by the simulator on the support of the prior of $\bm{\theta}$. That is, $q^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\bm{\Theta})\}$ (cannon2022investigating). For single observations, the definition of misspecification can be extended to using summary statistics (kelly2024misspecificationrobust) by defining $\bm{b}(\bm{\theta})=\mathbb{E}_{p(\bm{x}_{\rm sim}\mid\bm{\theta})}[\iota_{\omega}(\bm{x}_{\rm sim})]$ and $\bm{b}_{j}=\mathbb{E}_{q^{*}(\bm{x}_{\rm obs})}[\iota_{\omega}(\bm{x}_{\rm obs}^{(j)})]$, for each $\bm{x}_{\rm obs}^{(j)}\in O$ where $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ is the set of observations each associated to a different, unknown true $\bm{\theta^{(j)}}$ value. The simulator is said to be misspecified if there is no $\{\bm{\theta}\}_{i=1}^{N_{\rm obs}}$ in the support of the prior for which $\bm{b}(\bm{\theta}^{(j)})=\bm{b}_{j}$ for each $j$.
In this paper, we adopt the alternate, but similar, definition of misspecification provided in wehenkel2025addressing, which is more aligned with amortised SBI. They define a simulator to be misspecified if $\exists\mathcal{S}\subseteq\Theta\times\mathcal{X}_{\rm obs}:\forall(\bm{\theta},\bm{x}_{\rm obs})\in\mathcal{S}$
@@
This definition aligns with the amortised SBI task as it identifies misspecification as a set-wise phenomenon. Our goal is to recover robust and reliable posterior inference for all $\bm{x}_{\rm obs}\in\mathcal{X}_{\rm obs}$. This definition does, however, ignore situations when the prior distribution is misspecified. Prior misspecification is addressed at the end of the paper in the discussion.
@@ 2.4 Importance weighted autoencoders
Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, kingma2013auto), which introduces a latent space ${\Theta}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\xi}}(\bm{x}|\,\bm{\theta})$ and $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x})$, known as the decoder and the encoder respectively. The vanilla VAE aims to maximise the evidence of the data using the evidence lower bound
$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})}\left[\log p_{\bm{\xi}}(\mathbf{x}|\mathbf{\theta})\right]-\mathrm{KL}\!\left(q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})\,\|\,p(\mathbf{\theta})\right)$ (5)
with respect to $\bm{\xi}$ and $\bm{\phi}$, where $p(\bm{\theta})$ is a prior over the latent parameter. However, using this bound has been shown to induce mode-seeking behaviour, leading to overly simplified representations and poor inference. burda2015importance introduced IWAE, which is based on a strictly tighter evidence lower bound derived from importance sampling. The log-evidence for a single point is given as
$\log p_{\bm{\theta}}(\bm{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\xi}}\left(\textbf{x},\bm{\theta}^{(l)}\right)}{q_{\bm{\phi}}\left(\bm{\theta}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\xi},\bm{\phi};\textbf{x}\right)$ (6)
$\text{with}\quad\bm{\theta}^{(l)}\sim q_{\bm{\phi}}\left(\bm{\theta}|\textbf{x}\right),$ (7)
which is a mass-covering objective that targets the evidence. We use the The doubly reparameterized gradient (DReG) estimator (tucker2018doubly) to avoid low signal-to-noise in the gradient estimation.
cremer2017reinterpretingimportanceweightedautoencoders proposed using the implicit distribution given by importance sampling from the true posterior using the variational posterior as a proposal distribution. Given samples $\bm{\theta}_{2:N_{s}}:=\bm{\theta}_{2},...,\bm{\theta}_{N_{s}}$, they define the implicit distribution arising from importance sampling as
$\tilde{q}_{IW}(\mathbf{\theta}|\,\mathbf{x},\bm{\theta}_{2:N_{s}}):=\frac{p(\bm{x},\bm{\theta})}{\frac{1}{N_{s}}\Big(\frac{p(\bm{x},\bm{\theta})}{q_{\mathbf{\phi}}(\mathbf{\theta}|\,\mathbf{x})}+\sum_{k=1}^{N_{s}}\frac{p(\bm{x},\bm{\theta}_{j})}{q_{\mathbf{\phi}}(\mathbf{\theta}_{j}|\,\mathbf{x})}\Big)},$ (8)
which has the property that
$q_{\text{SIR}}(\bm{\theta}|\,\bm{x}):=\mathbb{E}_{\mathbf{\theta}_{2:N_{s}}}[\tilde{q}_{IW}(\mathbf{\theta}|\,\mathbf{x},\bm{\theta}_{2:N_{s}})]\rightarrow p(\mathbf{\theta}|\,\mathbf{x}),\,\rm{when}\,N_{s}\rightarrow\infty$ (9)
the expectation of the implicit distribution approaches the true posterior. Sampling from this distribution is detailed explicitly in cremer2017reinterpretingimportanceweightedautoencoders, and is done by sampling-importance-resampling (SIR).
@@ 3 Method: RVNP
Motivated by a desire to have a method that can recover robust amortised posterior inference with a fixed embedding space such that the simulation-to-reality gap is bridged with an interpretable, flexible, and data-driven error model, we extend neural likelihood estimation and the error modelling approaches for model misspecification (ward2022robust; kelly2024misspecificationrobust) to amortised SBI and propose robust variational neural posterior estimation (RVNP). RVNP builds on variational methods for solving the inverse problem under a learnt likelihood in SBI (glockler2022variational) by adopting an importance weighted autoencoder amortised variational inference scheme (burda2015importance) to define robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T).
In what follows, we drop the explicit embedding notation $\iota_{\omega}$ unless necessary. Our goal is to use a normalising flow parametrized by $\bm{\phi}$, $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$, that can approximate the true posterior distribution in an amortised manner. Algorithm 1 provides an overview of the RVNP and RVNP-T algorithms.
@@ 3.1 Generative model
Throughout, we assume a pre-training or fixing of the neural statistic embedding $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ occurred. The first step of RVNP is to train the normalising flow $p_{\bm{\Psi}}(\bm{x}|\,\bm{\theta})$ to approximate the likelihood using the generated samples from the simulator by minimising the NLE objective (Equation 3). Once the normalising flow has been trained, we assume that $p(\bm{x}_{\rm sim}|\,\bm{\theta})\approx p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ and include the surrogate in the forward model.
@@ 3.2 Variational posterior
$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]\,p(\bm{\theta}^{(l)})p(\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha}))}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]$ (12)
where $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, $\bm{\alpha}$ are the weights of the neural network $\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]$ using a Monte Carlo (MC) estimate. We use the logsumexp function to ensure the MC estimates are stable.
@@ 3.3 Posterior tuning
$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})}[\log q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (13)
This final objective can be identified with the noisy neural posterior estimation (NNPE, ward2022robust) objective and the error augmentation method suggested by Cranmer_2020. However, our error model does not have to be globally fixed and has been inferred using variational inference.
@@ 3.4 Error modelling
Theoretically, any model can be introduced within the RVNP framework. However, we opted to only include error models that can inflate the error on the simulator model. The default error model that we adopt in RVNP is given by the Gaussian covariance matrix
@@
a neural network that outputs the diagonal components of the covariance matrix and the non-diagonal components $\bm{\Lambda}$ are globally inferred. The second model we include is a global Gaussian model
@@
defined through a full rank Gaussian covariance matrix that is parametrised in terms of a Cholesky decomposition. This error model is constant across the parameter space and does not explicitly depend on $\bm{\theta}$.
Our method generalises very easily to include any inductive bias that we believe explains the simulation-to-reality gap. Our approach can be understood in terms of the simulator defining a population prior $p(\bm{\theta},\bm{x}_{\rm sim})=p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})p(\bm{\theta})$ and the error model representing the likelihood of $\bm{x}_{\rm sim}$ for the observed data.
@@ 3.5 Prior over the correction model parameters
In RVNP, a (pseudo-)prior over the error model parameters $p(\bm{\xi}(\bm{\theta}))$ can be introduced to regularise the influence of the error model and make the inference more robust in the single-posterior case. In our experiments, we only use a uniform prior distribution over the error model parameters, as we focus on the robustness imparted by multiple observations.
Algorithm 1 Robust Variational Neural Posterior Estimation (RVNP)
[…]
2:repeat
3:  Sample minibatch $B\subset D$
4:  $\mathcal{L}_{\rm NLE}(\bm{\Psi})\leftarrow-\mathbb{E}_{B}[\log p_{\bm{\Psi}}(\bm{x}_{\rm sim}\mid\bm{\theta})]$
5:  $\bm{\Psi}\leftarrow\bm{\Psi}-\eta\nabla_{\bm{\Psi}}\mathcal{L}_{\rm NLE}$
6:until convergence
7:return $p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta})$
8:
9:LearnPosteriorAndErrorModel($p_{\bm{\Psi}^{*}}$, $\mathcal{O}=\{\bm{x}_{\rm obs}\}$)
10:repeat
11:  for each $\bm{x}_{\rm obs}\in$ minibatch $\mathcal{O}$ do
12:   $\{\bm{\theta}^{(l)}\}_{l=1}^{K}\sim q_{\bm{\phi}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
13:   for $l=1$ to $K$ do
14:    $\{\bm{x}_{\rm sim}^{(l,m)}\}_{m=1}^{M}\sim p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta}^{(l)})$
15:    $w^{(l)}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\bm{\xi}_{\alpha}(\bm{\theta}^{(l)})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim}^{(l,m)})p(\bm{\theta}^{(l)})p(\bm{\xi}_{\alpha})}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\bm{x}_{\rm obs})}$
16:   end for
17:   $\mathcal{L}_{V}\leftarrow-\log\left(\frac{1}{K}\sum_{l=1}^{K}w^{(l)}\right)$
18:   $(\bm{\phi},\bm{\alpha})\leftarrow(\bm{\phi},\bm{\alpha})-\eta\nabla_{(\bm{\phi},\bm{\alpha})}\mathcal{L}_{V}$
19:  end for
20:until convergence
21:return $q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$, $p_{\bm{\xi}_{\alpha}^{*}(\bm{\theta})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim})$
22:
23:TunePosterior($q_{\bm{\phi}^{*}}$, $p_{\bm{\xi}_{\alpha}^{*}}$, $D$) // Optional: RVNP-T
24:repeat
25:  Sample $(\bm{x}_{\rm sim},\bm{\theta})\sim D$, $\bm{x}_{\rm obs}\sim p_{\bm{\xi}_{\alpha}^{*}(\bm{\theta})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim})$
26:  $\mathcal{L}_{\rm tune}(\bm{\phi})\leftarrow-\log q_{\bm{\phi}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
27:  $\bm{\phi}\leftarrow\bm{\phi}-\eta\nabla_{\bm{\phi}}\mathcal{L}_{\rm tune}$
28:until convergence
29:return $q_{\bm{\phi}^{**}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
30:
31:Infer($\bm{x}_{\rm obs}$, $N_{\rm samples}$) // With optional SIR
32:for $i=1$ to $N_{\rm samples}$ do
33:  $\bm{\theta}^{(i)}\sim q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
34:end for
35: Optionally apply sample-importance-resampling to $\{\bm{\theta}^{(i)}\}$
36:return $\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm samples}}$